I cant seem to find the answer to this math problem... let along try to solve it.
the question is: if f(x) = 3x^2(squared) then find f(4) and f(3b)

Answer by praseenakos@yahoo.com(507)

QUESTION:

if f(x) = 3x^2(squared) then find f(4) and f(3b)

ANSWER:

if f(x) = 3x^2(squared) then find f(4) and f(3b)

f(x) = 3 ( x )^2


To find f(4), replace x by 4 in the given equation as follows.

==> f(4) = 3 ( 4 )^2

==> = 3 * 16

==> = 48


To find f(3b), replace x by 3b in the given equation as follows.


f(3b) = 3 ( 3b )^2


= 3 * 9 b^2

= 27 b^2

If the second question is f(3),

f(3) = 3 ( 3 )^2


= 3 * 9


= 27

Hope you understood.

Regards.

praseenakos@yahoo.co.in